40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 19 18:10:10 2022, Paco, Update, BHD, Part V of BHR upgrade - POX11 path, LO path, and ITMX Oplev DJI_0167.JPGDJI_0169.JPGDJI_0168.JPG
    Reply  Wed Apr 20 08:20:22 2022, Paco, Update, BHD, Part V of BHR upgrade - PR2 weirdness 
       Reply  Wed Apr 20 18:50:03 2022, Koji, Update, BHD, Part V of BHR upgrade - PR2 weirdness Screenshot_2022-04-20_18-46-11.png
    Reply  Thu Apr 21 16:49:01 2022, Paco, Update, BHD, Part V of BHR upgrade - BS Oplev + LO1 offset 
       Reply  Thu Apr 21 18:42:47 2022, Tega, Update, BHD, Part V of BHR upgrade - Align LO and AS beams to BHD BS 
          Reply  Thu Apr 21 20:33:31 2022, Koji, Update, BHD, AS1 UR OSEM problem localized in the chamber Screenshot_from_2022-04-21_20-13-13.png
             Reply  Fri Apr 22 12:03:09 2022, Tega, Update, BHD, AS1 UR OSEM problem localized in the chamber 
                Reply  Fri Apr 22 12:15:08 2022, Anchal, Update, BHD, Cable post installation 
                   Reply  Mon Apr 25 16:57:57 2022, Anchal, Update, BHD, Cable post installation 
Message ID: 16797     Entry time: Thu Apr 21 16:49:01 2022     In reply to: 16788     Reply to this: 16800
Author: Paco 
Type: Update 
Category: BHD 
Subject: Part V of BHR upgrade - BS Oplev + LO1 offset 

[Paco, JC, Yuta]

We aligned the BS oplev using the new BSOL mirror pair. The main change is now the AOI of the oplev on the BS is quite normal. The output beam in the in-air table was quite large (diverging?) so we had to place a short FL lens in front of the QPD.


Separately, I added the LO1 YAW offset of ~ -2500 counts (before the coil driver changes it was -24500 counts) and saw LO beam hitting LO2. This means the alignment of the LO beam can move downstream.

ELOG V3.1.3-